f(x) = 2x -1 g(x) = Square root (x-1) Find the domain of f, g, f+g, f-g, fg, ff, f/g, and g/f

f(x) = 2x -1
g(x) = Square root (x-1)
Find the domain of f, g, f+g, f-g, fg, ff, f/g, and g/f.

I actually got the answers to the domain of f and g, but i'm confused on how to find the others based on just f and g.
Like for the domain of f, I got (-infinity, positive infinity)
then for the domain of g, (negative infinity, 1) U (1, positive infinity)
Answer #1
(x) = 2x -1 g(x) = Square root (x-1) Find the domain: f: all real numbers g: x>=1 f+g: x>=1 because sqrt(x-1) is one of the terms f-g: x>=1 becaue sqrt (x-1) is one of the terms fg= f(sqrt(x-1))= 2sqrt(x-1)-1: x>=1 same reason ff: all Real numbers f/g = (2x-1)/sqrt(x-1) Domain is x > 1 g/f = sqrt(x-1)/(2x-1) Domain is x>=1 ===================== Cheers, Stan H.
answered by: Help me please :P